If the gravitational field strength at the top of Mount Everest is 9.772 N/kg, approximately how tall (in




feet) is the mountain?

1
Expert's answer
2022-06-13T08:22:50-0400
"g(h)=G\frac{M}{(R+h)^2}=G\frac{M}{R^2}*\frac{R^2}{(R+h)^2}"

"g(h)=g(0)\frac{R^2}{(R+h)^2}"

"9.772=9.807*\left(\frac{6378}{6378+h}\right)^2"

"\frac{6378}{6378+h}=0.998"

"h=11.41\:\rm km=37 434\: ft"